Suppose x has a distribution with μ = 40 and σ = 19.

Find P(36 ≤ x ≤ 41).
Answer #1
for normal distribution z score =(X-μ)/σx
here mean=       μ= 40
std deviation   =σ= 19.000
probability =P(36<X<41)=P((36-40)/19)<Z<(41-40)/19)=P(-0.21<Z<0.05)=0.5199-0.4168=0.1031